Calculate the equilibrium constant for the reaction N2(g) + 3H2(g) - 2NH3(g) at 25°c given that ∆G =32.90kjmol-1.
1
Expert's answer
2020-11-10T14:04:37-0500

"\Delta" G= -RTlnk

R=universal gas constant=8.314 Jule/mole K,

T=Kelvin temperature=298K,"\Delta" G=32.90*1000=32900Jule

Putting all these value in above equation we gat

K=1.72*10^-6